DX LUCI/AO Commissioning

01:00 System powered on and ready for preset

01:02 Preset, AO800, MODE=ACQUIRE, GS=0

01:04 No Star so David point and collimate

01:13 Preset, AO800, MODE=ACE, gs=11, rs=0
  • Guide star found
  • Ref star seen

01:17 RunAO
  • On center pupil CCD39 went Not Live
    • Switch fastlink fiber and slope arrive
  • Optimize Gain 1.7, 1.6, 0.6 (too high)
    • Get a few skip frames
    • Set to 1.15, 0.7, 0.6
  • RunAO successful

01:28 Gopt not running
  • 012928
  • help, (ee->control())->c2m()
  • Stop/start gopt data and gopt processes
  • Gopt work

01:41 Skip Frame * Open loop from AOS GUI * AdSec Arb die

01:44 Preset 400 modes (good seeing table)
  • RunAO
  • Fix commitReadout Problem

Bin 1, 990 Hz, 400 Modes

  • Note, background image not used: not correctly collected. SR probably not correct (higher?)

02:03 Data Sets
  • Bin 1, 990 Hz, Gopt=0.23
  • Guider FWHM = 0.8, DIMM not working
  • FeII, Clear
  • 020300, 020352, 020429
  • SR = 37, 30, 32
TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20171209_020300 ??? 9.0 245734 14 NaN 1.07 b1m400 1 400 990 0.3 0.1 0.1 3 150 ON 37.3 FeII clear 157 10 ONSKY zero.fits  
20171209_020352 ??? 9.0 245068 13 NaN 1.07 b1m400 1 400 990 0.3 0.1 0.1 3 151 ON 30.9 FeII clear 157 10 ONSKY zero.fits  
20171209_020429 ??? 9.0 244601 9 NaN 1.16 b1m400 1 400 990 0.3 0.1 0.1 3 150 ON 31.8 FeII clear 157 10 ONSKY zero.fits  
  • 020300 PSF:
    020300_psf.jpg
  • 020300 modal:
    020300_modal.jpg
  • 020300 Jitter:
    020300_jitter.jpg

02:09
  • DIMM = 1.3
  • OH1060, Z
  • SR may be in error or PSF center problem
  • 020957, 021028, 021043
  • SR = ??

Bin 1, 990 Hz, 300 Modes

  • Note, background image not used: not correctly collected. SR probably not correct (higher?)

02:15 Preset 300 modes (bad Seeing Table)
  • 1.4, 0.7, 0.7 (manual)
  • 022535, 022611, 022641
  • SR = 52, 42, 47
  • Rotator at 286
TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20171209_022535 ??? 9.1 228601 14 1.24 1.04 b1m300 1 300 990 0.4 0.2 -1.0 3 142 ON 52.4 FeII clear 157 10 ONSKY zero.fits  
20171209_022611 ??? 9.1 228140 11 1.25 1.27 b1m300 1 300 990 0.4 0.2 -1.0 3 139 ON 41.6 FeII clear 157 10 ONSKY zero.fits  
20171209_022641 ??? 9.1 227761 13 1.25 1.02 b1m300 1 300 990 0.4 0.2 -1.0 3 144 ON 47.1 FeII clear 157 10 ONSKY zero.fits  
  • 022535 PSF:
    022535_psf.jpg
  • 022535 Modal:
    022535_modal.jpg
  • 022535 Jitter:
    022535_jitter.jpg

Bin 2, 625 Hz, 156 Modes

  • Note, background image not used: not correctly collected. SR probably not correct (higher?)

02:33 Preset AO1064, bin 2, 11.2 mag sent
  • see 12.2 seen
  • 625 Hz
  • 2.25, 0.52, 0.44 (auto)
  • set to 1.75, 0.72, 0.59
  • Gopt = 0.15
  • NCPA on

02:57 Had to open/close loop to stop/start wfsarb because was not processing LUCI images

02:59
  • rot angle 113
  • 025932, 025959, 030016
  • SR = 37, center problem, 37
TrackNo RefStar Mag El Wind DIMM/OL Rec bin #mod freq gain mod nph AntiDrift SR filter exp #frames disturb SN notes
20171209_025932 ??? 12.2 241761 12 1.19 1.10 b2m153 2 153 625 0.4 0.2 0.1 3 48 ON 37.4 FeII clear 157 10 ONSKY zero.fits  
20171209_030016 ??? 12.2 241220 10 1.10 1.18 b2m153 2 153 625 0.4 0.2 0.1 3 47 ON 37.3 FeII clear 157 10 ONSKY zero.fits  
  • 025932 Modal:
    025932_modal.jpg
  • 025932 Jitter:
    025932_jitter.jpg

03:05 Preset to bin 3 star, 200 Hz
  • test new pupil center routine
  • 2.0, 1.0, 1.0
  • rot angle 133
  • no telemetry collected

03:27 Preset to AO807, bin 4
  • 200 Hz
  • 3.0, 2.0, 2.0
  • no telemetry
  • pupil centering seems to work fine.

03:45 Hand over to Jochen for LUCI/AO commissioning

04:01 First Com target
  • Bin 2, 990 Hz
  • 2.0, 0.6, 0.5
  • Gopt=0.22

04:07 Loop Open, skip frame
  • used "close loop" button

04:14 Jochen reported motion of star on slit. AO TT LUT is not on. Will turn it on for next object

05:02 Preset
  • bin 2, 990 Hz
  • 1.97, 0.6, 0.5
  • gopt=0.27
  • NCPA on
  • TT LUT on
  • Seeing 0.9

06:31 Preset to same star with different PA
  • 2.0, 0.6, 0.6
  • gopt=0.24
  • NCPA on
  • TT LUT on

07:16 Loop Open, skip frame, Gopt goes to 0.03

07:29 Loop Closed again. Looked like gains were overdriving

07:39 Finally finished new preset

09:11 Preset to 11.5 mag star

09:18 RunAO failed -> Seeing from 1.7" to 3.3"

-- %USERSIG{DougMiller - 2017-12-09}%
I Attachment Action Size Date Who Comment
020300_jitter.jpgjpg 020300_jitter.jpg manage 112 K 09 Dec 2017 - 05:56 DougMiller 020300 Jitter
020300_modal.jpgjpg 020300_modal.jpg manage 89 K 09 Dec 2017 - 05:47 DougMiller 020300 modal
020300_psf.jpgjpg 020300_psf.jpg manage 86 K 09 Dec 2017 - 05:36 DougMiller 020300 PSF
022535_jitter.jpgjpg 022535_jitter.jpg manage 119 K 09 Dec 2017 - 05:57 DougMiller 022535 Jitter
022535_modal.jpgjpg 022535_modal.jpg manage 98 K 09 Dec 2017 - 05:57 DougMiller 022535 Modal
022535_psf.jpgjpg 022535_psf.jpg manage 82 K 09 Dec 2017 - 05:57 DougMiller 022535 PSF
025932_jitter.jpgjpg 025932_jitter.jpg manage 120 K 09 Dec 2017 - 06:15 DougMiller 025932 Jitter
025932_modal.jpgjpg 025932_modal.jpg manage 89 K 09 Dec 2017 - 06:14 DougMiller 025932 Modal
Topic revision: r4 - 09 Dec 2017, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback